Marcus drives from one town to another in 6 hours. on the return trip, his speed is increased by 10 mph and the trip takes 5 hours. Find his rate on the return trip. how far apart are the towns?
Answer by nyc_function(2741)

6x = 5(x + 10)
6x = 5x + 50
6x - 5x = 50
x = 50
His going rate = x or 50mph.
His returning rate = 50 + 16 or 60 mph.
The towns are 300 miles apart.